OpenStudy (anonymous):

What is the specific heat of an unknown substance if 2000 J of energy are required to raise the temperature of 4 grams of the substance 5 degrees Celsius? Specific heat= (Joules)/(celsius) (grams)

OpenStudy (anonymous):

e = mst where s = specific heat 2000 = 4 * s * 5 s = 100

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

Just plug in those numbers and solve for s. 2000 / 20 = 100